OpenStudy (anonymous):

Could someone help me, please? 1. 2n > 14 n > _____ (1 point) 2. n/5≥ 11 n ≥ _______ (1 point) 3. –3y ≤ –18 y ≥ _______ (1 point)

OpenStudy (anonymous):

You'll need to isolate n, 1. Divide both sides by 2 2. Multiply both sides by 5. 3.Divide both side by -3

OpenStudy (anonymous):

1) \[2n > 14\] Divide both sides by 2 n > 7 2)\[\frac{ n }{ 5 }\ge 11\] multiply both sides by 5 \[n \ge 55\] 3)\[-3y \le -18\] Divide both sides by -3. Side you divide by a negative, flip the sign. \[y \ge 6\]

OpenStudy (anonymous):

Thank you both, for explaining this! Greatly appreciated! I'm kind of new to "open study". How do I give medals??

OpenStudy (anonymous):

You click on the blue button that says "Best Response" and welcome :)